# A Low-Rank ADMM Splitting Approach for Semidefinite Programming

### 14 Mar 2024 v1

@@ 4.3 The Dynamic Rank Selection Strategy
To improve the approach’s performance, one challenge lies in how to apply a smaller rank while maintaining the algorithm’s stability. The majority of literature on low-rank SDP algorithms commonly advocates for employing a rank of $\sqrt{2m}$ to maintain favorable properties of local minima. However, theoretical advancement by So et al. (2008) suggests that an approximate solution of satisfactory quality can be efficiently achieved by employing a logarithmic scale rank of $m$. The theorem underscoring this idea is stated as follows:
@@
This theoretical insight encourages us to adventurously implement a rank of $O(\log(m))$ within the approach to be proposed.
Although a lower rank reduces storage requirements and computational costs, employing a rank $r$ that does not satisfy $r\geq\sqrt{2m}$ may render the problem challenging to solve. In certain cases, consistently employing a rank of order $O(\log(m))$ may result in slow problem-solving or even failure to solve the problem. To mitigate this, our approach is straightforward: start with a very small rank (of order $O(\log(m))$) and keep the rank as low as possible, increasing it only when necessary. We found it is good enough to apply the rank adjustment only in the warm-start phase since it is empirically shown that LRADMM is more capable of handling low-rank instances than LRALM. To determine when to increase the rank, we devise a heuristic method to detect when a problem becomes "difficult" to solve, where the number of inner iterations of the ALM subproblems reflects the difficulty. Upon detecting the difficulty, we interrupt the current subproblem’s iterative cycle, augment the rank by a multiplier, $\alpha$, and then proceed with the algorithm’s iterations. The results presented in Table 2 compare the performance of LoRADS when applying a consistent rank of $r=2\log(m)$, $r=\sqrt{2m}$, or employing the dynamic rank selection strategy, where the rank is multiplied by a factor of 1.5 each time. We continue to use the set of selected representative problems in the previous subsection to illustrate the concept, with the implementation and experiment settings following the descriptions in Section 5.1 and Section 5.2. From Table 2, it can be concluded that in certain cases, using a rank of $2\log(m)$ significantly reduces solving times compared to $\sqrt{2m}$. However, for some other problems, a rank of $2\log(m)$ makes the problem slower to solve (e.g., qap7) or even unsolvable (e.g., G64, qap7, qap10, theta12). However, the dynamic rank strategy consistently capitalizes on the benefits of low rank while mitigating the difficulty introduced.
@@ 4.4 Solving the ADMM Subproblems with the Warm-started CG
Since we consider the linear SDP problems, the subproblems (11) and (12) of LRADMM can be reformulated into the following linear systems with dimension $nr\times nr$:
@@
To address the linear systems discussed, we employ a warm-started CG method. In this approach, we initialize the algorithm with $V^{k}$ (or $U^{k}$) to efficiently solve the linear system for $U^{k+1}$ (or $V^{k+1}$). It is important to highlight that the inclusion of the $\gamma$ penalty term from the splitting significantly improves the conditioning of the linear systems (22) and (23). Consequently, when $\gamma$ is sufficiently large, these systems are well-conditioned and can be effectively solved using the CG method without any preconditioning. To demonstrate the efficacy of this method, we present in Table 3 the number of ADMM iterations, the total CG iterations required to solve the problem, and the average number of CG iterations per ADMM iteration (involving solving two subproblems) across the selection of representative problems. The problems are solved solely using LRADMM with out Burer-Monteiro warm-start and the ranks are set to be $\sqrt{2m}$. The result clearly shows that the average number of CG iterations per subproblem is extremely lower than the worst-case scenario of $nr$ iterations.
@@ 5 Numerical Experiments
In this section, we provide a detailed empirical evaluation of LoRADS applied to a diverse set of SDP problems. We commence by outlining the key implementation details and the experimental setup, providing a clear context for the conducted experiments in Section 5.1. Subsequently, in Section 5.2, we assess the performance of our method on three sets of SDP problems: MaxCut, Matrix Completion, and a collection of general SDP benchmark problems.
@@ 5.1 Implementation Details
We implement LoRAS in ANSI C with open-source numerical linear algebra libraries, lapack (Anderson et al., 1999) and blas (Blackford et al., 2002). To solve the subproblems of LRADMM, we use the warm-started conjugate gradient method to solve the linear systems (22) and (23), where $f(UV^{\top})=\big{\langle}C,UV^{\top}\big{\rangle}$. For the dynamic rank adjustment strategy discussed previously, we initiate the rank at $\log(2m)$ and increment the rank by a factor of 1.5 with each rank update. Regarding the timing for variable splitting, we switch to Phase II upon meeting a switching tolerance of 1e-3 in the majority of instances tested. However, for all MaxCut problems (identified by starting with ’G’ and ’mcp’), we apply a more lenient switching tolerance of 1e-2. Our code is capable of automatically selecting the phase-switching tolerance by analyzing the structural characteristics of the problem. The error used as the indicator for terminating Phase I is primal infeasibility, which is consistent with the final termination indicator outlined in Section 5.2. Beyond these specific settings, our investigation suggests that the algorithm’s effectiveness relies on the application of certain techniques and implementation strategies. In the ensuing subsection, we delve into these aspects in detail.
@@
The efficacy of the ADMM is significantly influenced by the choice of the penalty parameter, $\rho$. A well-known method for adaptively selecting this parameter is the residual balancing technique, introduced by He et al. (2000); Boyd et al. (2011), whose idea is to balance primal and dual residuals. However, our experiments indicate that this method does not consistently provide satisfactory performance in our algorithm, primarily due to the challenge of selecting appropriate values for the related hyper-parameters across different problem instances. To avoid the need for problem-specific manual adjustment of the penalty parameter, we use a simple heuristic strategy that incrementally increases the penalty parameter within our approach. Specifically, with in LRADMM, we increase $\rho$ by $1.2$ times every five iterations until it reaches a predetermined maximum value, $\rho_{max}$, which is set to be $5000$ in all the experiments. The initial value of $\rho$ for LRADMM inherits from the warm start phase. In the warm-start phase, the update strategy for the update strategy for $\rho$ follows that described by Burer and Choi (2006).
@@
One consideration is that the appropriate value of $\rho$ for the ADMM and the ALM may differ when transitioning between phases. To address this, we introduce a heuristic factor, $h$, to adjust rho appropriately during phase transitions. Specifically, when switching from LRALM to LRADMM, we multiply $\rho$ by the heuristic factor $h$ instead of carrying it over directly. Overall, the directly inheriting $\rho$ from the warm-start phase is proved to be robust enough in the majority of cases. However, we have observed that in certain types of problems, selecting an appropriate heuristic factor $h$ can significantly improve the convergence performance. In subsequent experiments, we applied the heuristic factor as follows: $h=10$ for all MaxCut, MaxCut MinBisection, graph partition, and quadratic assignment problems [^2] The problems are recognized by: ”G” and ”mcp” for MaxCut; G40_mb and G48_mb for MaxCut MinBisection; ”gpp” for graph partition; and ”qap” for quadratic assignment problems.; in matrix completion problems, we set $h=5$ for MC_1000 to MC_8000, and $h=2.5$ for MC_10000 to MC_40000; for all other problems, we set $h=1$, indicating no use of the heuristic factor.
@@
To improve the performance of the LRALM during the warm-start phase, we incorporate enhancements suggested in Burer and Choi (2006). However, we do not apply the rank updating method introduced in their work. In general, our implementation follows that of SDPLR v1.03333For the source code, see https://yalmip.github.io/solver/sdplr/. For a brief guide, see https://sburer.github.io/files/SDPLR-1.03-beta-usrguide.pdf.
@@ 5.2 Experiment Setting
To benchmark our method, we compared its performance with the following widely used SDP solvers:
@@
  • SDPLR (Burer and Choi, 2006). An open-sourced solver based on the Burer-Monteiro low-rank factorization method. The comparison can show the power of the low-rank and matrix splitting techniques in our method.
@@
  • SDPNAL+ (Sun et al., 2019). Open-sourced solver based on ALM, aiming at solving large-scale SDP problems.
@@
  • COPT (Ge et al., 2022). Commercial solver based on interior point method. It is currently the best performing SDP solver in Hans Mittelmann’s benchmarks for optimization software on SDP problems [^4] https://plato.asu.edu/ftp/sparse_sdp.html..
    The solver serves as an interior point method benchmark.
@@
Given that different solvers employ distinct termination settings, some of which are rigidly embedded in their code, we have made efforts to match their termination criteria while acknowledging differences. These solvers terminate when certain errors meet a given tolerance $\epsilon$:
@@
  • LoRADS and SDPLR terminate when $\frac{\left\|\mathcal{A}\left(X\right)-b\right\|}{1+\left\|b\right\|_{\infty}}\leq\epsilon$.
@@
  • SDNAL+ terminates when $\max\left\{\frac{\|\mathcal{A}(X)-b\|}{1+\|b\|},\frac{\|\mathcal{A}^{*}(y)-C\|}{1+\|C\|}\right\}<\epsilon,$ where $\mathcal{A}^{*}(y)=\sum_{i=1}^{m}y_{i}A_{i}.$
    In the original work by Sun et al. (2019), two additional errors are defined. However, in our setting, both of them are zeros.
@@
  • COPT terminates when $\max\left\{\frac{\|\mathcal{A}(X)-b\|}{1+\|b\|_{1}},\frac{|\min\{0,\sigma_{\min}(C-\mathcal{A}^{*}(\lambda))\}|}{1+\|c\|_{1}},\frac{\langle C,X\rangle-\lambda^{\top}b}{1+|\langle C,X\rangle|+|\lambda^{\top}b|}\right\}<\epsilon,$ where $c$ is the flatten vector of matrix $C$.
We heuristically consider to aligned the termination criteria of LoRADS with that of SDPLR, since the second term in (13), $\frac{\gamma\|V^{k+1}-V^{k}\|_{F}}{1+\|V^{k}\|_{F}}$, is typically good when the first term is met, namely the primal infeasibility.
This alignment allows us to directly evaluate the perform differences between LoRADS and SDPLR. Although the termination criterion is solely based on primal infeasibility, it is important to note that the measurement of primal infeasibility is normalized using the infinity norm $\|b\|_{\infty}$. This approach is more strict compared to utilizing the $\|b\|$ or $\|b\|_{1}$, particularly in instances that have numerous constraints.
Since each solver terminates with distinct criteria, we present a unified report of three commonly defined errors for the solutions of each solver in Appendix B.
Primal Infeasibility:
$\frac{\|\mathcal{A}(X)-b\|}{1+\|b\|_{1}},$
Dual Infeasibility:
$\frac{|\min\{0,\sigma_{\min}(C-\mathcal{A}_{X}^{*}(\bm{\lambda}))\}|}{1+\|c\|_{1}},$
Primal Dual Gap:
$\frac{\langle C,X\rangle-\bm{\lambda}^{\top}b}{1+|\langle C,X\rangle|+|\bm{\lambda}^{\top}b|}.$
The three error measures are chosen for ranking in Hans Mittelmann’s benchmarks, and it is also widely used in many existing solvers, like COPT (Ge et al., 2022) and HDSDP (Gao et al., 2022).
@@
In the following experiments in this section, we set $\epsilon$ to be $51e-$ 5, and the time limit is 10000 seconds [^5] The instances failed to be solved in 10000 seconds are marked with ”t”. All the experiments are performed on a Macbook Pro with a 12 Core M3 Pro chip and 18GB RAM.
@@ 5.3 The MaxCut Problems
Given a weighted graph $G$ with weight $w$, the MaxCut is a famous problem with seeks to divide the vertices of a graph into two disjoint subsets, maximizing the sum of weights of the edges connecting vertices in different subsets. We can formulate the SDP form of the MaxCut problem as follows:
@@
where $L(G,w)$ is the Laplacian matrix is defined by
@@
We test all the previously mentioned solvers on the Gset [^6] For problem files, see https://web.stanford.edu/~yyye/yyye/Gset/ (Ye, 2003). As shown in Table 4, LoRADS notably outperforms the other solvers in almost all cases. Initial observations indicate that SDPNAL+ is less effective in solving this particular family of SDP problems. Remarkably, the advantages of LoRADS become more pronounced as the scale of the problem increases. For problems with $n,m<5000$, LoRADS is approximately $1$ to $2$ times faster than both SDPLR and COPT. However, for larger instances where $n,m\geq 5000$, the speed advantage of LoRADS increases to around $10$ to $20$ times faster than SDPLR and COPT. Such observations underscore that the solving times of LoRADS increase much less rapidly with the scale of the problem compared to SDPLR and COPT, highlighting the efficiency of employing the low-rank ADMM approach.
@@
G1 800 0.60 1.35 50.95 0.69 G35 2000 1.56 5.65 871 3.05
G2 800 0.58 1.52 99.17 0.71 G36 2000 2.50 9.70 1051 2.77
G3 800 0.59 2.08 52.13 0.93 G37 2000 1.70 6.39 1117 2.64
G4 800 0.66 1.70 38.33 0.73 G38 2000 1.51 9.02 1146 2.75
G5 800 0.60 1.76 70.21 0.62 G39 2000 2.56 6.41 854 3.14
G6 800 0.49 1.64 118 0.80 G40 2000 2.73 8.97 856 3.18
G7 800 0.47 1.22 34.96 0.55 G41 2000 2.68 7.16 866 3.31
G8 800 0.45 1.15 32.67 0.72 G42 2000 2.72 7.37 792 3.07
G9 800 0.45 1.16 97.88 0.58 G43 1000 0.30 0.70 67.02 0.76
G10 800 0.48 1.22 48.58 0.78 G44 1000 0.31 0.68 41.62 0.61
G11 800 0.07 0.59 294 0.23 G45 1000 0.32 0.86 40.61 0.93
G12 800 0.06 0.30 223 0.28 G46 1000 0.31 0.77 66.14 0.62
G13 800 0.05 0.22 229 0.23 G47 1000 0.30 1.12 104 1.07
G14 800 0.29 0.97 134 0.58 G48 3000 0.14 0.61 5221 2.22
G15 800 0.34 0.79 134 0.54 G49 3000 0.16 0.81 2516 2.11
G16 800 0.31 0.76 141 0.59 G50 3000 0.20 0.84 4928 2.07
G17 800 0.30 0.77 121 0.53 G51 1000 0.50 1.19 134 0.79
G18 800 0.36 0.82 132 0.38 G52 1000 0.55 1.28 140 0.68
G19 800 0.42 0.94 155 0.48 G53 1000 0.42 2.91 126 0.73
G20 800 0.27 0.59 142 0.47 G54 1000 0.50 1.33 133 0.54
G21 800 0.32 0.75 164 0.45 G55 5000 1.17 5.86 t 18.62
G22 2000 0.80 4.48 438 3.39 G56 5000 0.97 5.74 t 17.66
G23 2000 0.88 4.31 447 3.49 G57 5000 0.79 10.46 t 11.64
G24 2000 0.80 5.08 435 3.60 G58 5000 20.65 87.35 t 23.85
G25 2000 0.74 5.40 440 3.56 G59 5000 20.32 109 t 26.70
G26 2000 0.93 3.35 499 3.62 G60 7000 1.70 10.23 t 40.00
G27 2000 1.61 2.45 434 3.28 G61 7000 1.44 15.94 t 41.34
G28 2000 1.95 3.84 453 4.03 G62 7000 1.52 22.86 t 23.57
G29 2000 0.84 3.61 441 4.52 G63 7000 34.01 190 t 58.38
G30 2000 1.78 2.98 421 4.71 G64 7000 45.39 260 t 64.07
G31 2000 1.87 3.45 430 3.86 G65 8000 1.79 32.35 t 34.70
G32 2000 0.27 1.92 2904 1.64 G66 9000 2.09 39.91 t 48.33
G33 2000 0.24 1.74 2003 1.65 G67 10000 2.38 53.78 t 59.26
G34 2000 0.23 1.33 3241 1.67
5.4 The Matrix Completion Problems

### 25 Mar 2024 v2

@@ 4.3 The Dynamic Rank Selection Strategy
To improve the computation and memory efficiency, it is often desirable to apply a smaller rank while maintaining the algorithm’s stability. Most literature on low-rank SDP algorithms typically recommends a rank of $\sqrt{2m}$ to maintain favorable properties of local minima. However, a theoretical result by So et al. (2008) suggests that an approximate solution of satisfactory quality can be efficiently achieved by employing a rank of $O(\log m)$. The theorem underscoring this idea is stated as follows:
@@
This theoretical insight motivates us to aggressively adopt an $O(\log(m))$ rank to further reduce storage requirements and computational costs. Nevertheless, employing a rank $r$ that does not satisfy $r\geq\sqrt{2m}$ may make the problem more challenging. In some scenarios, fixing a rank of $O(\log(m))$ may result in slow convergence or even failure to find a solution. Our strategy is simple: start with a small rank order (of $O(\log(m))$), increasing it only when necessary. We observed that adjusting the rank solely during the warm-start phase suffices, as LRADMM is empirically more capable of handling low-rank instances than LRALM. To determine when to increase the rank, we employ a heuristic to identify when a problem becomes "difficult" to solve, reflected by the number of inner iterations of the ALM subproblems. Upon detecting the difficulty, we halt the current subproblem’s iterative cycle, increase the rank by a factor $\alpha$, and then proceed with the algorithm’s iterations. The results presented in Table 2 compare the performance of LoRADS when applying a consistent rank of $r=2\log(m)$, $r=\sqrt{2m}$, or employing the dynamic rank selection strategy, where the rank is multiplied by a factor of 1.5 each time. We continue to use the selected problems in the previous subsection to illustrate the concept, with the experiment settings following the descriptions in Section 5.1 and Section 5.2. From Table 2, it can be seen that using a rank of $2\log(m)$ significantly reduces solving times compared to $\sqrt{2m}$ in many cases, while, for some other cases, a rank of $2\log(m)$ makes the problem slower to solve (e.g., qap7) or even unsolvable (e.g., qap10, theta12). On the other hand, the dynamic rank strategy can leverage the advantages of a low-rank structure while bypassing its potential challenges.
@@ 4.4 Solving the ADMM Subproblems with the Warm-started CG
Since we consider the linear SDP problems, where the subproblems (11) and (12) of LRADMM with respect to $(U^{k+1},V^{k+1})$ can be reformulated into the following linear systems with dimension $nr\times nr$:
@@
To address the linear systems discussed, we employ a warm-started CG method. In this approach, we initialize the algorithm with $V^{k}$ (or $U^{k}$) to efficiently solve the linear system for $U^{k+1}$ (or $V^{k+1}$). It is important to highlight that the inclusion of the $\gamma$ penalty term from the splitting significantly improves the conditioning of the linear systems (22) and (23). Consequently, when $\gamma$ is sufficiently large, these systems are well-conditioned and can be effectively solved using the CG method without any preconditioning. To demonstrate the efficacy of this method, we present in Table 3 the number of ADMM iterations, the total CG iterations required to solve the problem, and the average number of CG iterations per ADMM iteration (involving two subproblems) across the selection of representative problems. The problems are solved solely using LRADMM without Burer-Monteiro warm-start, and the ranks are set to be $\sqrt{2m}$. The result clearly shows that the average number of CG iterations per subproblem is extremely smaller than the worst-case scenario of $nr$ iterations.
@@ 5 Numerical Experiments
In this section, we provide a detailed empirical evaluation of LoRADS applied to a diverse set of SDP problems. We begin by outlining the key implementation details and the experiment setup in Section 5.1. Subsequently, in Section 5.2, we assess the performance of our method on three sets of SDP problems: MaxCut, Matrix Completion, and a collection of general SDP benchmark problems.
@@ 5.1 Implementation Details
We implement LoRADS [^2] LoRADS is avaliable at https://github.com/COPT-Public/LoRADS in ANSI C with open-source numerical linear algebra libraries, Lapack (Anderson et al., 1999) and Blas (Blackford et al., 2002). To solve the subproblems of LRADMM, we use the warm-started conjugate gradient method to solve the linear systems (22) and (23), where $f(UV^{\top})=\big{\langle}C,UV^{\top}\big{\rangle}$. For the dynamic rank adjustment strategy discussed previously, we initiate the rank at $2\log(m)$ and increment the rank by a factor of 1.5 with each rank update. Specially, given the extensive scale of matrix completion problems from MC_10000 to MC_40000, a modified approach is adopted where the initial rank is set at $\log(m)$. Regarding the timing for variable splitting, we switch to Phase II upon meeting a switching tolerance of 1e-3 in the majority of instances tested. However, for all MaxCut problems (all problems in Section 5.3 and problems identified by starting with ’mcp’ in Section 5.5), we apply a more lenient switching tolerance of 1e-2. Our code is capable of automatically selecting the phase-switching tolerance by analyzing the structural characteristics of the problem. The error used as the indicator for terminating Phase I is primal infeasibility, consistent with the final termination indicator outlined in Section 5.2. Beyond these specific settings, our investigation suggests that the algorithm’s effectiveness relies on applying certain techniques and implementation strategies. In the ensuing subsection, we delve into these aspects in detail.
@@
The efficacy of ADMM is significantly influenced by the choice of the penalty parameter, $\rho$. A well-known adaptive method for selecting this parameter is the residual balancing technique, introduced by He et al. (2000); Boyd et al. (2011), whose idea is to balance primal and dual residuals. However, our experiments indicate that this method does not consistently provide satisfactory performance, primarily due to the challenge of selecting appropriate values for the related hyper-parameters across different problem instances. To avoid the need for problem-specific manual adjustment of the penalty parameter, we use a simple heuristic strategy that incrementally increases the penalty parameter within our approach. Specifically, within LRADMM, we increase $\rho$ by $1.2$ times every five iterations until it reaches a predetermined maximum value, $\rho_{\max}=5,000$. The initial value of $\rho$ for LRADMM inherits from the warm start phase. In the warm-start phase, the update strategy for $\rho$ follows from Burer and Choi (2006).
@@
The appropriate value of $\rho$ for ADMM and ALM may differ when transitioning between phases. To address this, we introduce a heuristic factor, $h$, to adjust $\rho$ appropriately during phase transitions. When switching from LRALM to LRADMM, we multiply $\rho$ by $h$ instead of carrying it over directly. In most cases, directly inheriting $\rho$ from the warm-start phase exhibits sufficiently robust performance. However, for certain types of problems, selecting an appropriate heuristic factor $h$ can significantly improve the performance. In subsequent experiments, we applied the heuristic factor as follows: $h=10$ for all MaxCut, MaxCut MinBisection, graph partition, and quadratic assignment problems [^3] The problems are recognized by: ”G” and ”mcp” for MaxCut; G40_mb and G48_mb for MaxCut MinBisection; ”gpp” for graph partition; and ”qap” for quadratic assignment problems.; in matrix completion problems, we set $h=5$ for MC_1000 to MC_8000, and $h=2.5$ for MC_10000 to MC_40000; for all other problems, we set $h=1$ to indicate no use of the heuristic factor.
@@
To improve the performance of the LRALM during the warm-start phase, we incorporate enhancements suggested in Burer and Choi (2006). However, we do not apply the rank-updating method introduced in their work. Our implementation generally follows that of SDPLR v1.03444For the source code, see https://yalmip.github.io/solver/sdplr/. For a brief guide, see https://sburer.github.io/files/SDPLR-1.03-beta-usrguide.pdf..
@@ 5.2 Experiment Setting
We compare LoRADS with the following widely used SDP solvers:
@@
  • SDPLR (Burer and Choi, 2006). An open-source solver that utilizes the Burer-Monteiro (BM) low-rank factorization method.
@@
  • SDPNAL+ (Sun et al., 2019). An open-source solver employing ALM, designed for tackling large-scale SDP problems.
@@
  • COPT (Ge et al., 2022). A commercial solver based on the interior point method. COPT is currently recognized as the top-performing solver in Hans Mittelmann’s benchmarks for SDP problems [^5] https://plato.asu.edu/ftp/sparse_sdp.html.. This solver is included as a benchmark for the interior point method.
@@
Given the diverse termination settings employed by different solvers—some of which are deeply integrated into their code—we have tried aligning their termination criteria while recognizing inherent differences. These solvers terminate when certain error metrics fall below a specified tolerance $\epsilon$:
@@
  • LoRADS and SDPLR terminate when $\frac{\left\|\mathcal{A}\left(X\right)-b\right\|_{2}}{1+\left\|b\right\|_{\infty}}\leq\epsilon$.
@@
  • SDNAL+ terminates when $\max\left\{\frac{\|\mathcal{A}(X)-b\|_{2}}{1+\|b\|_{2}},\frac{\|\mathcal{A}^{*}(y)-C\|_{F}}{1+\|C\|_{F}}\right\}<\epsilon,$ where $\mathcal{A}^{*}(y)=\sum_{i=1}^{m}y_{i}A_{i}.$ Note that Sun et al. (2019) define two additional errors. However, both of them are zeros in our setting.
@@
  • COPT terminates when $\max\left\{\frac{\|\mathcal{A}(X)-b\|_{2}}{1+\|b\|_{1}},\frac{|\min\{0,\sigma_{\min}(C-\mathcal{A}^{*}(\lambda))\}|}{1+\|c\|_{1}},\frac{\langle C,X\rangle-\lambda^{\top}b}{1+|\langle C,X\rangle|+|\lambda^{\top}b|}\right\}<\epsilon,$ where $c$ is the flatten vector of matrix $C$.
We have chosen to align the termination criterion of LoRADS with that of SDPLR rather than directly employing the condition (13). This decision is informed by the observation that the second term in (13), $\frac{\gamma\|V^{k+1}-V^{k}\|_{F}}{1+\|V^{k}\|F}$, typically meets the required standards once the primal infeasibility gap is addressed. Notably, the primal infeasibility is measured as a value normalized using the infinity norm $\|b\|_{\infty}$, making this approach significantly more rigorous than utilizing either $\|b\|_{2}$ or $\|b\|_{1}$, particularly in scenarios involving a large number of constraints.
@@
In the following experiments in this section, we set $\epsilon$ to be 1e-5, and the time limit to be 10,000 seconds [^6] The instances failed to be solved in 10,000 seconds are marked with ”t”. The iteration cap for the LRADMM algorithm is set at 5,000. All the experiments are performed on a Macbook Pro with a 12 Core M3 Pro chip and 18 GB RAM. For clarity, this section only reports the results of solving times for each experiment. A report of properly defined and aligned solving errors for all solvers in all the experiments can be found in Appendix B.
@@ 5.3 The MaxCut Problems
Given a weighted graph $G$ with weights $w$, the MaxCut problem aims to partition the vertices into two disjoint subsets such that the total weight of the edges connecting vertices across the subsets is maximized. The SDP formulation of the MaxCut problem can be expressed as follows:
@@
where the Laplacian matrix $L(G,w)$ is defined by
@@
We evaluated all previously discussed solvers using the Gset benchmark [^7] For problem files, see https://web.stanford.edu/~yyye/yyye/Gset/ (Ye, 2003), and report the test results in Table 4. Our findings suggest that SDPNAL+ struggles with this specific set of SDP challenges, and LoRADS outperforms the other solvers in almost all cases. For problems with $n,m<5000$, LoRADS is approximately $1$ to $2$ times faster than both SDPLR and COPT. However, for larger instances where $n,m\geq 5000$, the performance gap widens significantly, with LoRADS being $10$ to $20$ times faster than SDPLR and COPT. This trend highlights that the advantages of LoRADS become more pronounced in larger-scale problems.
@@
G1 800 0.18 0.75 33.43 0.82 G35 2000 0.23 1.60 723 3.49
G2 800 0.17 0.68 36.49 0.78 G36 2000 0.24 1.41 828 2.99
G3 800 0.18 0.81 31.69 0.90 G37 2000 0.25 1.84 381 3.52
G4 800 0.17 0.69 29.59 0.81 G38 2000 0.25 1.85 414 3.25
G5 800 0.15 0.66 25.41 0.72 G39 2000 0.38 1.76 352 3.29
G6 800 0.23 0.67 30.18 0.61 G40 2000 0.57 2.82 428 3.54
G7 800 0.22 0.60 29.54 0.56 G41 2000 0.52 1.99 444 3.20
G8 800 0.20 0.69 34.49 0.61 G42 2000 0.54 1.67 360 3.19
G9 800 0.20 0.54 28.22 0.54 G43 1000 0.09 0.44 18.20 0.97
G10 800 0.20 0.64 36.14 0.81 G44 1000 0.11 0.37 18.17 0.74
G11 800 0.05 0.47 49.13 0.25 G45 1000 0.10 0.40 53.57 1.00
G12 800 0.05 0.34 58.34 0.26 G46 1000 0.10 0.56 50.16 0.85
G13 800 0.05 0.17 58.37 0.28 G47 1000 0.10 0.43 55.18 1.10
G14 800 0.07 0.28 45.07 0.50 G48 3000 0.12 0.67 1436 2.09
G15 800 0.07 0.34 55.99 0.59 G49 3000 0.14 0.76 462 2.00
G16 800 0.07 0.33 56.73 0.58 G50 3000 0.13 0.96 1433 1.88
G17 800 0.06 0.34 39.46 0.53 G51 1000 0.09 0.45 51.64 0.83
G18 800 0.12 0.37 30.17 0.37 G52 1000 0.09 0.38 58.41 0.88
G19 800 0.12 0.41 33.84 0.54 G53 1000 0.09 0.43 19.48 0.75
G20 800 0.12 0.32 34.12 0.48 G54 1000 0.08 0.48 49.06 0.66
G21 800 0.11 0.35 27.20 0.50 G55 5000 0.38 3.45 5587 18.04
G22 2000 0.28 1.66 277 3.42 G56 5000 0.42 4.03 4664 18.18
G23 2000 0.31 1.31 483 3.67 G57 5000 0.54 6.55 7123 11.34
G24 2000 0.27 1.91 274 3.50 G58 5000 0.87 14.76 8426 24.52
G25 2000 0.27 1.26 276 3.26 G59 5000 2.42 26.48 7318 25.16
G26 2000 0.27 1.72 286 3.76 G60 7000 0.66 7.66 t 40.20
G27 2000 0.40 1.66 525 3.49 G61 7000 0.86 6.69 t 41.16
G28 2000 0.50 1.99 418 4.09 G62 7000 1.02 16.25 t 23.59
G29 2000 0.33 2.37 272 4.27 G63 7000 1.92 23.69 t 59.32
G30 2000 0.31 1.37 553 4.54 G64 7000 5.60 61.82 t 62.78
G31 2000 0.41 2.03 321 3.86 G65 8000 1.30 24.80 t 33.35
G32 2000 0.18 1.36 645 1.50 G66 9000 1.44 29.39 t 43.09
G33 2000 0.17 1.21 475 1.44 G67 10000 1.58 42.73 t 57.18
G34 2000 0.17 1.17 579 1.61
5.4 The Matrix Completion Problem